# Universidad Andres Bello, Facultad de Ciencias Exactas, Departamento de Ciencias Físicas - Instituto de Astrofísica, Autopista Concepcion-Talcahuano 7100, Talcahuano, Chile Instituto de Física, Universidad de Antioquia, Calle 70 52-21, Medellín, Colombia

### v1

@@ 3.1 Atmospheric parameters
The atmospheric parameters of our targets were obtained through an interactive procedure based on the GAIA G, Gbp, Grp and 2MASS J, H, Ks photometry of the cluster. During this procedure, the effective temperature (T) and gravity (log g) of each RGB star (including our targets) were obtained, and at the same time, CMDs were corrected for differential reddening. First, we fitted a PARSEC isochrone (Bressan et al., 2012) to the RGB and Red Clump (RC), assuming an age of 13.0 Gyrs. We accounted for reddening by applying the Cardelli et al. (1989) reddening law to the isochrone. The visual absorption AV, the RV parameter, the intrinsic distance modulus (m-M)0 and the global metallicity [M/H] were determined by simultaneously fitting the RGB, RC and RGB-tip in the Ks vs. Gbp - Ks, G vs. Gbp - Grp, Gbp vs. Gbp - Grp and Ks vs. J - Ks CMDs. The T and log g of each RGB star were then determined as those corresponding to the point on the isochrone where the magnitude matches that of the star (we avoided using RC stars for this step). We used the Ks magnitude because it is the least affected by reddening and, consequently, by differential reddening. Having the temperature we obtained the intrinsic Gbp - Ks color of each star from the color-temperature relation of the RGB part of the isochrone and, by subtracting the mean reddening obtained from the isochrone fitting, also the differential reddening at the position of each star. Finally, for each star we selected the 4 closest neighbours (5 stars in total) and corrected its G, Gbp, Grp, J, H, Ks magnitudes using the mean differential reddening of the 5 stars. We used the Gbp - Ks color because it is the most sensitive to any reddening variation. This procedure was iterated until any improvement of the CMDs was negligible.
We obtained (m-M)0 = 15.32 0.05, AV = 1.91 0.05 and RV = 2.7 0.1, resulting in E(B-V) = 0.71, close to the Deras et al. (2023) value of 0.64 but substantially higher than the Harris (1996, 2010 edition) value of 0.54. This value is higher than any other from the literature since the RV we estimated (2.7) is lower than the value usually assumed (RV = 3.1). We also obtained [M/H] = 0.6 0.05, a value higher than the iron content of the cluster ([Fe/H] = 0.9) given by Deras et al. (2023) (and corroborated by us). This is not surprising since globular clusters are usually -enhanced. Also, the isochrone fitting gave an age estimation of 12 Gyrs, however, we cannot strongly rely on this estimate since no turn off point (TO) is visible in the Gaia CMDs. However, the deep HST data of Deras et al. (2023) do indeed reveal the TO and an age of 13.1 0.5 Gyr.
@@ 3.2 Chemical abundances
The method used to derive the elemental abundances was a local thermodynamics equilibrium (LTE) analysis using all the 10 targets with the Brussels Automatic Code for Characterizing High accuracy Spectra: BACCHUS (Masseron, Merle, & Hawkins, 2016)), which relies on the radiative code Turbospectrum (Alvarez & Plez, 1998; Plez, 2012) and the MARCS model atmosphere grid Gustafsson et al. (2008). With the effective temperature and surface gravity photometrically derived and an initial metallicity estimate from Deras et al. (2023) we allowed BACCHUS to determine and the convolution. The is obtained by minimizing the trend of Fe abundances against their reduced equivalent width, while the convolution parameter accounts for the total effect of instrument resolution, macroturbulence, and on the line broadening (Hawkins et al., 2016).
@@
Figure 2 shows the fit obtained for Oxygen, Aluminum and Nickel for the star 2M171639032807212 as an example of the quality of the fit described. In the Figure, the black dotted line is the observed spectrum of the target and the red continuum line is the created synthetic spectrum.
@@ 4.1 Results and comparison with other in-situ clusters
We report abundances for C, N, O, Mg, Al, Si, P, K, Ca, Ti, V, Cr, Mn, Fe, Ni and Ce. These are shown in Table 2 together with overall [/Fe] abundances for each star determined with the available elements (O, Mg, Si and Ca), weighted equally. Each abundance has its associated weighted error, determined by the square root of the quadratic sum of each individual error and dividing the result by the number of stars. Table 2 also contains the microturbulence velocities obtained by BACCHUS. This is the first spectroscopic report of abundances manually determined for this cluster and, since we used high resolution spectroscopy, the most robust study to the date. We obtain a mean metallicity for the cluster of [Fe/H] = 0.87 0.02, with no evidence of any intrinsic Fe abundance variation. Our metallicity is in very good agreement with that obtained by Deras et al. (2023), [Fe/H] = , and Conroy et al. (2018), [Fe/H] = . However, it is some 0.5 dex lower than the canonical value of 0.45 from the Harris (1996, 2010 edition) catalog and similarly offset from the Vásquez et al. (2018) value of 0.37. Our determination involves a large sample of stars observed at high resolution and SNR, and thus should be preferred to previous estimates.
@@
With our new metallicity, NGC 6316 is now no longer a member of the classic metal-rich BGCs. Indeed, based on both the Bica et al. (2024) and Geisler et al. (in prep) metallicity distribution functions of BGCs, which (assuming that it is indeed a BGC) are bimodal with peaks near -0.5 and -1.1, NGC 6316 falls in the lower metallicity domain, which is now known to be the dominant mode.
@@
Given our good agreement with the Deras et al. (2023) estimate, we note that this gives increased support to their age determination of 13.1 +/- 0.5 Gyr, ensuring that NGC 6316 is indeed a rather old in situ GC for its metallicity. It now falls nicely along the age-metallicity sequence of in situ GCs, which are substantially more metal-rich than their accreted cousins of similar age, and indeed older than similar metallicity ex situ GCs (Cohen et al., 2021; Belokurov & Kravtsov, 2024). If instead the old Harris (1996, 2010 edition) metallicity was used, NGC 6316 would be a substantial outlier in both of these diagrams.
@@
Figure 3 shows abundance relations for light elements involved in the MP phenomenon. Alongside NGC 6316 (white stars) there are four other clusters: NGC 6522 in emerald inverted triangles ([Fe/H] = , Fernández-Trincado et al. 2019), NGC 6558 in blue squares ([Fe/H] = , González-Díaz et al. 2023), Ton 2 in red triangles ([Fe/H] = , Fernández-Trincado et al. 2021) and NGC 6121 (M4) in Fuchsia circles ([Fe/H] = , Mészáros et al. 2020. These clusters were selected because all of their abundances were determined using BACCHUS, and they present similar metallicities as our analyzed stars. These criteria were selected in an attempt to find similar (or different) behaviours at similar/different location for in-situ GCs. According to Geisler et al. (in prep), all 5 of these GCs are in-situ, NGC 6316, NGC 6522 and NGC 6558 are indeed from the Bulge, Ton 2 is from the Disk, and it is uncertain whether NGC 6121 is from the Bulge or Disk.
@@
As seen in Figure 3 top left panel, NGC 6316 presents a strong C-N anti-correlation, as do NGC 6522, NGC 6558, Ton 2 and NGC 6121, but our GC presents the strongest and most extended anticorrelation of all the compared B/D clusters. Fernández-Trincado et al. (2021) report a correlation between Ce and N in Tonantzintla 2. We observe a similar correlation, with substantial scatter, in our cluster (bottom right panel), suggesting that this cluster may have received contamination from external sources. There is a significant spread in Al and Mg, with the latter spread being higher than any of the BGC compared in the plots. This spread in both elements is expected, as Carretta et al. (2009) establishes that Al-rich and Mg-depleted stars are present in clusters that are massive (M , quite metal-poor ([Fe/H] ), or both. Our cluster, although it is not considered ”metal poor”, has a mass of according to the Baumgardt database, which is sufficient to be considered massive. However, the bottom left panel does not show a linear Mg-Al anticorrelation but rather a scatter in both elements. Mészáros et al. (2020) defined 2P as [Al/Fe] . However, this definition is not reliable for our case, since two of the three stars above this limit are the lowest in Nitrogen (hence deemed 1P) in our sample, using the definition of [N/Fe] from Geisler et al. (2021). It is worth noting that these two nitrogen-low stars are also the highest in magnesium in the sample (top right panel in Figure 2. Taken as a whole, NGC 6316 displays most of the usual symptoms of the MP disease, but this is the first time it has been diagnosed in detail.
@@
Figure 4 shows a violin diagram comparing C, N, O, Mg, Al, Si, Ca, Fe, and Ce in NGC 6316 with those of the same clusters as in Figure 3: NGC 6522, Ton 2, NGC 6558 and NGC 6121. This diagram illustrates the range for each element, represented by a vertical line with two small horizontal lines at the limits and a third one in-between showing the mean abundance of the element. The color width of each element across the vertical line indicates the number of stars with each abundance, revealing how many stars contribute to the spread in each element.
@@
Our cluster shows a higher spread than the other in situ GCs in most of the elements shown in the diagram, with the exception of Si. Additionally, there is generally good agreement among these GC abundances and spreads. M4 presents higher mean abundances than all the other GCs in O, Mg, Al, Si and Ce.
4.2 Comparisons with ASPCAP
Given the substantial number of new abundances obtained, it is a good opportunity to compare them with the ASPCAP values provided in the APOGEE value-added catalogue from Schiavon et al. (2024), but also taking into account that a more complete and detailed analysis is being carried out in Geisler et al. (in prep.) for all CAPOS clusters. Schiavon et al. (2024) presented abundances for a group of stars in NGC 6316, which includes all 10 of our target stars, enabling a reliable comparison. For brevity, we will refer to our results as BACCHUS and those from Schiavon et al. (2024) as ASPCAP.
The mean iron abundance in BACCHUS (0.87 0.02) is approximately 0.1 dex lower than ASPCAP (0.78 0.005), yet still aligns with the metal-poor side of the measurements when comparing with the old references. ASPCAP also presents a lower metallicity dispersion than BACCHUS. One of the most notable differences between the ASPCAP and BACCHUS abundances lies in Carbon. Although both methods show the same anticorrelation in C and N and similar trend, the two most Carbon-depleted stars, identified as 2P stars, are 0.3-0.4 dex more depleted in BACCHUS than in ASPCAP. For Nitrogen, while ASPCAP and BACCHUS exhibit similar dispersion, ASPCAP reports a mean nitrogen abundance that is 0.1 dex lower than BACCHUS. Regarding the two mentioned 2P stars, their mean nitrogen values are similar in both methods, but the BACCHUS abundances show significantly higher dispersion.
The -elements O, Mg, Si, and Ca show a difference of 0.05 in their overall [/Fe] abundance, with BACCHUS reporting a mean of [/Fe] = 0.30 0.02 (Figure 5), while ASPCAP reports a mean [/Fe] = 0.25. Notable differences arise in individual abundances. For instance, oxygen shows similar mean values between ASPCAP ([O/Fe] = 0.28) and BACCHUS ([O/Fe] = 0.26), but BACCHUS exhibits significantly higher dispersion (0.04 vs 0.18). Since MPs are present in this cluster, an oxygen dispersion is expected, lending more credibility to our individual oxygen abundances. Indeed, a C-O correlation (or N-O anticorrelation) is not observed using ASPCAP values.
Magnesium shows a strong agreement in its mean abundance between the two methods, but BACCHUS again exhibits higher dispersion. Notably, two stars in BACCHUS are more magnesium-depleted than any ASPCAP star. Among these, one also displays the high carbon depletion and enhanced nitrogen mentioned earlier, while the other presents a normal level of carbon but enhanced nitrogen abundance (and thus both considered as 2P). Aluminum and Potassium show excellent agreement in both their mean abundance and dispersion across methods. Both ASPCAP and BACCHUS report the same dispersion for Silicon, but BACCHUS abundances are 0.1 dex higher. Calcium in BACCHUS has a slightly higher abundance, with dispersion also being twice as large as in ASPCAP. Titanium shows similar dispersion in both methods but ASPCAP reports abundances that are 0.05 dex higher. However, the typical errors of our Ti measurements allow to determine an agreement between ASPCAP and BACCHUS for this element.
It should be noted that using our [Fe/H] and [/Fe] abundances to calculate the global metallicity [M/H] via the formula from Salaris, Chieffi, & Straniero (1993), we obtain [M/H] = -0.66 0.03. This result is very close to our photometric determination in section 3.1, further supporting the validity of our abundance and atmospheric parameter determinations. Conversely, the global metallicity reported by ASPCAP is almost identical to their iron metallicity, differing by only 0.01 dex, which does not align with their calculation using the aforementioned formula.
@@ 4.3 The in-situ nature of NGC6316
Bottom: Mg vs Al-planes but using BACCHUS abundances from Mészáros et al. (2020) and CAPOS papers. The abundances were systematically shifted as indicated in the appendix A of Belokurov & Kravtsov (2024). Left panel shows Circles representing single stars, while diamonds represents mean abundances for each cluster. Color dashed lines are the contours of the distribution of the single stars. The distribution of Mg and Al are shown in histograms at the right and top of the plot, respectively. Single star distributions show two clear dense areas where the in-situ and ex-situ stars are concentrated. This is more difficult to reproduce with mean values. The right panel shows individual stars of some CAPOS clusters compared in the previous section. The stars tend to cluster in an intermediate region between the two denser areas, slightly below the separation line in the ex-situ zone. Using only 1P stars, NGC 6316 and Ton 2 fall in the in-situ region.
As mentioned, Massari, Koppelman, & Helmi (2019) and Geisler et al. (in prep) classify NGC 6316 as an in-situ GC. Belokurov & Kravtsov (2024, hereafter BK24) proposed an abundance-based classification scheme for in-situ and ex-situ GCs, utilizing [Mg/Fe] vs [Al/Fe] abundances derived from ASPCAP data and other sources scaled to ASPCAP (see their Fig. 11, right panel). We replicated this scheme (Figure 6, top left panel) using eight GCs—four in-situ (NGC 6316, NGC 6388, NGC 6397, and Ton 2) and four ex-situ (NGC 288, NGC 1851, NGC 2808, and NGC 3201)—based on the classification given by Massari, Koppelman, & Helmi (2019), with [Mg/Fe] and [Al/Fe] abundances determined using BACCHUS (Mészáros et al. 2020 and related studies discussed earlier in this section) shifting the abundances as indicated in BK24 (see their Appendix A). Our analysis focused exclusively on 1P stars, as both Mg and Al elements are influenced by MPs. The selection of 1P stars was done trying to find an agreement with the [Al/Fe] = 0.3 limit suggested by Mészáros et al. (2020), with the lowest [N/Fe] and highest [C/Fe] stars, giving priority to the [N/Fe] and [C/Fe] levels given any contradiction with [Al/Fe]. The division line between in- and ex-situ proposed by BK24 was also included. NGC 6558 and NGC 6522 were excluded from this plot because, despite having at least one 1P star with [Al/Fe] ¡ 0.3, their [N/Fe] levels were too high. Using BACCHUS abundances, we found that the trend line proposed by BK24 did not function as expected. The separation between in-situ and ex-situ GCs was unclear, and although NGC 6316 appears in the ”in-situ” side of the plot, the separation among in- and ex-situ GCs seems to be more complicated than simply using a line.
BK24 also mentions that including 2P stars should not significantly affect the plot, as the anomalously low [Mg/Fe] abundances are counteracted by elevated [Al/Fe] levels. This shifts the GCs in the plot from the top-left to the bottom-right, parallel to the dividing line, while keeping them on the same side as when using only 1P stars. Based on this statement, we included both 1P and 2P stars from these GCs, as well as other GCs whose samples did not appear to contain 1P stars. (Figure 6, top right panel)
We found that although adding 2P stars does not drastically change the distribution of the majority of GCs in the plot, it does for NGC 6316, which exhibits MP characteristics (”MP disease”) M13 and NGC 6397, that change their positions to the other side. Thus, the addition of 2P stars is not reliable. Also, the dividing line proposed by BK24 continues to perform poorly when using 2P stars.
@@
Another analysis was conducted using all GCs from Mészáros et al. (2020) with more than five stars, systematically shifting the data following the BK24 methodology to align the [Mg/Fe] and [Al/Fe] abundances with the BK24 separation line distinguishing in-situ and ex-situ GCs (Figure 6, bottom left panel). In there, circles represent individual stars from Mészáros et al. (2020), while diamonds indicate the mean abundance for each cluster. The color-dashed lines outline the distribution contours of individual stars. The upper and right subplots display histograms of the overall dataset’s [Mg/Fe] and [Al/Fe] distributions. The single-star distributions exhibit two distinct dense regions, corresponding to the concentrations of in-situ and ex-situ stars. However, this distinction is less clear in the mean values (diamond symbols), where the two groups overlap within a broader area.
Although 2P stars are included in this plot, they are not critical in this context. Stars with non-pristine chemical compositions (2P stars) tend to scatter across the plot without impacting the dense regions formed by in-situ and ex-situ 1P stars. In addition to the two primary dense regions, a secondary, less prominent cluster of in-situ stars is observed at approximately [Al/Fe] 0.3, [Mg/Fe] 0.3, closer to the main concentration of ex-situ stars. The BK24 separation line crosses through the two main distributions but judging by eye, it seems that the slope should be more vertical with BACCHUS abundances.
This plot largely reproduces BK24’s main result: in-situ and ex-situ stars can be distinguished in the [Mg/Fe] vs. [Al/Fe] diagram. However, the separation is not entirely clean.
The right panel displays the same contours and dashed lines as in the left panel. Star symbols represent stars from the CAPOS clusters compared in Figures 3 and 4. After applying the shift proposed by BK24, these stars tend to cluster in an intermediate region between the two densest areas, slightly below the BK24 separation line in the ex-situ zone. When considering the mean with all 1P and 2P stars, all GCs fall into the ex-situ region. However, when using only 1P stars, NGC 6316 and Tonantzintla 2 are located in the in-situ zone, while NGC 6558 and NGC 6522, whose stars all lie in the ex-situ zone, lack 1P stars in the sample, agreeing with, or at least not contradicting, the classification of Massari, Koppelman, & Helmi (2019) and Geisler et al. (in prep).
Although these results align with the existing literature, it is important to note a methodological contradiction: Our analysis works best (albeit with substantial confusion between classes), only while using 1P stars, discarding the reliability of using 2P stars, while the separation line defined by BK24 was derived from both 1P and 2P stars.
A more detailed study focusing on 1P stars in all CAPOS clusters is currently being conducted by Geisler et al. (in prep.).

### v2

@@ 3.1 Atmospheric parameters
The atmospheric parameters of our targets were obtained through an interactive procedure based on the GAIA DR3 G, Gbp, Grp and 2MASS J, H, Ks photometry of the cluster. During this procedure, the effective temperature (T) and gravity (log g) of each RGB star (including our targets) were obtained, and at the same time, CMDs were corrected for differential reddening. First, we fitted a PARSEC isochrone (Bressan et al., 2012) to the RGB and Red Clump (RC), assuming an age of 13.0 Gyrs. We accounted for reddening by applying the Cardelli et al. (1989) reddening law to the isochrone. The visual absorption AV, the RV parameter, the intrinsic distance modulus (m-M)0 and the global metallicity [M/H] were determined by simultaneously fitting the RGB, RC and RGB-tip in the Ks vs. Gbp - Ks, G vs. Gbp - Grp, Gbp vs. Gbp - Grp and Ks vs. J - Ks CMDs. The T and log g of each RGB star were then determined as those corresponding to the point on the isochrone where the magnitude matches that of the star (we avoided using RC stars for this step). We used the Ks magnitude because it is the least affected by reddening and, consequently, by differential reddening. Having the temperature we obtained the intrinsic Gbp - Ks colour of each star from the colour-temperature relation of the RGB part of the isochrone and, by subtracting the mean reddening obtained from the isochrone fitting, also the differential reddening at the position of each star. Finally, for each star we selected the 4 closest neighbours (5 stars in total) and corrected its G, Gbp, Grp, J, H, Ks magnitudes using the mean differential reddening of the 5 stars. We used the Gbp - Ks colour because it is the most sensitive to any reddening variation. This procedure was iterated until any improvement of the CMDs was negligible.
We obtained (m-M)0 = 15.32 0.05, AV = 1.91 0.05 and RV = 2.7 0.1, resulting in E(B-V) = 0.71, close to the D23 value of 0.64 but substantially higher than the 0.56 value obtained by VFO07, and the H10 value of 0.54. Indeed, the value is an average between Zinn (1985) and Reed, Hesser, & Shawl (1988), who obtained E(B-V) values of 0.47 and 0.61, respectively, with the latter supporting a higher reddening. Our reddening is higher than any other from the literature for this cluster since we estimated RV = 2.7, lower than the value usually assumed (RV = 3.1) but in excellent agreement with Nataf et al. (2013), who determined RV 2.5 with a dispersion 0.2 toward the inner Galaxy. We also obtained [M/H] = 0.6 0.05, a value higher than the iron content of the cluster ([Fe/H] = 0.9) given by D23 (and corroborated by us). This is not surprising since globular clusters are usually -enhanced. Also, the isochrone fitting gave an age estimation of 12 Gyrs, however, we cannot strongly rely on this estimate since no turn off point (TO) is visible in the Gaia CMDs. However, the deep HST data of D23 do indeed reveal the TO and an age of 13.1 0.5 Gyr.
@@ 3.2 Chemical abundances
The method used to derive the elemental abundances was a local thermodynamics equilibrium (LTE) analysis using all the 10 targets with the Brussels Automatic Code for Characterizing High accuracy Spectra: BACCHUS (Masseron, Merle, & Hawkins, 2016)), which relies on the radiative code Turbospectrum (Alvarez & Plez, 1998; Plez, 2012) and the MARCS model atmosphere grid Gustafsson et al. (2008). With the effective temperature and surface gravity photometrically derived and an initial metallicity estimate from D23 we allowed BACCHUS to determine and the convolution. The is obtained by minimizing the trend of Fe abundances against their reduced equivalent width, while the convolution parameter accounts for the total effect of instrument resolution, macro-turbulence, and on the line broadening (Hawkins et al., 2016).
@@
Figure 2 shows the fit obtained for Oxygen, Aluminium and Nickel for the star 2M171639032807212 as an example of the quality of the fit described. In the Figure, the black dotted line is the observed spectrum of the target and the red continuum line is the created synthetic spectrum.
@@ 4.1 Results and comparison with other in-situ clusters
We report abundances for C, N, O, Mg, Al, Si, P, K, Ca, Ti, V, Cr, Mn, Fe, Ni and Ce. These are shown in Table 2 together with overall [/Fe] abundances for each star determined with the elements Si and Ca, weighted equally. Each abundance has its associated weighted error, determined by the square root of the quadratic sum of each individual error and dividing the result by the number of stars. We adopted the solar reference abundances from Asplund, Grevesse, & Sauval (2005) for all elements, as used in most CAPOS papers, due to their widespread adoption and consistency in the literature. Table 2 also contains the micro-turbulence velocities obtained by BACCHUS. This is the first spectroscopic report of abundances manually determined for this cluster and, since we used high resolution spectroscopy, we believe the most robust study to the date. We obtain a mean metallicity for the cluster of [Fe/H] = 0.87 0.02, with no evidence of any intrinsic Fe abundance variation. Our metallicity is in very good agreement with that obtained by D23, [Fe/H] = , and Conroy et al. (2018), [Fe/H] = . However, it is some 0.5 dex lower than the canonical value of 0.45 from the H10 catalogue and similarly offset from the Vásquez et al. (2018) values of , and . As mentioned, these values were obtained using the calcium triplet method. The metallicity reported in H10 is a weighted average of the values derived by AZ88 (scaled to C09), VFO07 and C09 (triple-weighted). The AZ88 value was obtained by averaging their calcium triplet [Fe/H] estimate with that from ZW84. C09, in turn, recalibrated the ZW84 scale to derive their own metallicities. ZW84 itself is based on a transformation of the Q39 index, derived primarily from the CaII K lines, among other indicators. Thus, the high metallicities reported for this cluster by these studies likely reflect issues related to the calcium-based methods. Our metallicity determination should therefore be preferred over previous estimates since it is also based on a large sample of stars observed at high resolution and SNR. This is further supported by the fact that our study - along with that of D23 - is one of the only investigations conducted exclusively on this cluster, providing a level of detail and accuracy that is not achievable when working on large samples of GCs.
@@
With our new metallicity, NGC 6316 is now no longer a member of the classic metal-rich BGCs. Indeed, based on both the Bica et al. (2024) and Geisler et al. (2025) metallicity distribution functions of BGCs, which (assuming that it is indeed a BGC) are bimodal with peaks near -0.5 and -1.1, NGC 6316 falls in the lower metallicity domain, which is now known to be the dominant mode.
@@
Given our good agreement with the D23 estimate, we note that this gives increased support to their age determination of 13.1 0.5 Gyr, ensuring that NGC 6316 is indeed a rather old in situ GC for its metallicity. It now falls nicely along the age-metallicity sequence of in situ GCs, which are substantially more metal-rich than their accreted cousins of similar age, and indeed older than similar metallicity ex situ GCs (Cohen et al., 2021; Belokurov & Kravtsov, 2024). If instead the old H10 metallicity was used, NGC 6316 would be a substantial outlier in both of these diagrams.
We obtained an overall [/Fe] = 0.31 0.02 determined using our silicon and calcium abundances. Oxygen and magnesium were excluded from the [/Fe] calculation, as these elements can vary substantially in globular clusters due to their involvement in MPs. To assess the reliability of our results, we determined the global metallicity [M/H] of the cluster using an alternative approach. Based on our derived [Fe/H] and [/Fe] abundances, we obtained an overall metallicity of [M/H] using the formula from Salaris, Chieffi, & Straniero (1993), in very good agreement with the value obtained through isochrone fitting ([M/H]).
@@
Figure 3 shows abundance relations for light elements involved in the MP phenomenon. Alongside NGC 6316 (white stars) there are four other clusters: NGC 6522 in emerald inverted triangles ([Fe/H] = , Fernández-Trincado et al. 2019), NGC 6558 in blue squares ([Fe/H] = , González-Díaz et al. 2023), Ton 2 in red triangles ([Fe/H] = , Fernández-Trincado et al. 2021) and NGC 6121 (M4) in Fuchsia circles ([Fe/H] = , Mészáros et al. 2020. These clusters were selected because all of their abundances were determined using BACCHUS, and they present similar metallicities as our analysed stars. These criteria were selected in an attempt to find similar (or different) behaviours at similar/different location for in-situ GCs. According to Geisler et al. (2025), all 5 of these GCs are in-situ, NGC 6316, NGC 6522 and NGC 6558 are indeed from the Bulge, Ton 2 is from the Disk, and it is uncertain whether NGC 6121 is from the Bulge or Disk.
@@
As seen in Figure 3 top left panel, NGC 6316 presents a strong C-N anti-correlation, as do NGC 6522, NGC 6558, Ton 2 and NGC 6121, but our GC presents the strongest and most extended anti-correlation of all the compared B/D clusters. There is a significant spread in Al and Mg, with the latter spread being almost equal to M4 and higher than the other GCs compared in the plots. This spread in both elements is expected, as Carretta et al. (2009b) establishes that Al-rich and Mg-depleted stars are present in clusters that are massive (M ), quite metal-poor ([Fe/H] ), or both. Our cluster, although it is not considered ”metal poor”, has a mass of according to the Baumgardt database, which is sufficient to be considered massive. However, the bottom left panel does not show a linear Mg-Al anti-correlation but rather a scatter in both elements. Mészáros et al. (2020) defined 2P as [Al/Fe] . However, this definition is not reliable for our case, since two of the three stars above this limit are the lowest in Nitrogen (hence deemed 1P) in our sample, using the definition of [N/Fe] from Geisler et al. (2021). It is worth noting that these two nitrogen-low stars are also the highest in magnesium in the sample (bottom left panel in Figure 3). Taken as a whole, NGC 6316 displays most of the usual symptoms of the MP disease, but this is the first time it has been diagnosed in detail.
@@
Figure 4 shows a violin diagram comparing C, N, O, Mg, Al, Si, Ca, Fe, and Ce in NGC 6316 with those of the same clusters as in Figure 3: NGC 6522, Ton 2, NGC 6558 and NGC 6121. This diagram illustrates the range for each element, represented by a vertical line with two small horizontal lines at the limits and a third one in-between showing the mean abundance of the element. The colour width of each element across the vertical line indicates the number of stars with each abundance, revealing how many stars contribute to the spread (the difference among the star with the highest abundance and that with the least abundance) seen in each element.
@@
Our cluster shows a higher spread than the other in situ GCs in C, N and Ce, with only M4 surpassing it in O, Mg and Ca. Additionally, there is generally good agreement among these GC abundances and spreads, with the exception of N and Al.
Despite knowing that ASPCAP does not perform optimally for stars with extreme types of “non-standard” elemental abundance patterns, we took the opportunity to compare our abundances with the ASPCAP values provided in the APOGEE value-added catalogue from Schiavon et al. (2024). This comparison is available in the appendix A.
@@ 4.2 The in-situ nature of NGC 6316
As mentioned, Massari, Koppelman, & Helmi (2019) and Geisler et al. (2025) classify NGC 6316 as an in-situ GC. Belokurov & Kravtsov (2024, hereafter BK24) proposed an abundance-based classification scheme for in-situ and ex-situ GCs, utilizing [Mg/Fe] vs [Al/Fe] abundances derived from ASPCAP data and other sources scaled to ASPCAP (see their Fig. 11, right panel). We replicated this scheme using all GCs from Mészáros et al. (2020) with more than five stars, systematically shifting the data following the BK24 methodology (see their Appendix A) to align the [Mg/Fe] and [Al/Fe] abundances with the BK24 separation line distinguishing in-situ and ex-situ GCs (Figure 5). In there, semi-transparent circles represent individual stars from Mészáros et al. (2020), while diamonds indicate the mean abundance for each cluster. The colour-dashed lines outline the distribution contours of individual stars. The upper and right subplots display histograms of the overall dataset’s [Mg/Fe] and [Al/Fe] distributions. The single-star distributions exhibit two distinct dense regions, corresponding to the concentrations of in-situ and ex-situ stars. However, this distinction is less clear in the mean values (diamond symbols), where the two groups overlap within a broader area.
We found that the method proposed by BK24 did not function as expected. The separation among in- and ex-situ GCs seems to be more complicated than simply using a line.
However, given the concentration of in- and ex- situ individual stars, it seems that a possible division can be made using a subgroup of stars, most probably 1P stars, that do not suffer from abundance variations. This subject, will be further explored in greater detail in a forthcoming paper (Geisler et al., 2025) focusing on 1P stars from all CAPOS clusters.